A isosceles triangle ABC has A= 45 degree angle and b=c=3 what is the length of a squared ?
Answer by amarjeeth123(569)

Using cosine rule we have a^2=b^2+c^2-2bccosA
Plugging in the values we get,
a^2=9+9-2(3)(3)cos45
a^2=18-18(0.7071)
a^2=5.272
Answer=5.272